OpenStudy (anonymous):

Help please 10 questions I already did 5 please correct me if they are wrong.

OpenStudy (anonymous):

Here are the questions @divu.mkr @madrockz @Opcode please @ someone i will be gone for a couple hours thanks :)@radar

OpenStudy (anonymous):

@hartnn can you help me please

OpenStudy (anonymous):

u want us to check ur answers?or to answer those ques which r not done..

OpenStudy (anonymous):

yea check the answers I did and you can do the other ones corect me if im wrong @madrockz

OpenStudy (anonymous):

@madrockz can yuo help?

OpenStudy (anonymous):

let me try.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

@RadEn can you help me please?

OpenStudy (raden):

1. T 4. T 5. F 8. F 9. F

OpenStudy (anonymous):

@RadEn what does that mean T And F mean

OpenStudy (raden):

T = true, F =false

OpenStudy (anonymous):

can you correct me with the ones i did wrong?

OpenStudy (anonymous):

like tell me if they are a b c or d so i can understand them?

OpenStudy (raden):

4. the factor of g^2 - 28g + 196 = (g - 14)(g - 14) it has the same factor, its g-14

OpenStudy (raden):

5. use the difference square formula : a^2 - b^2 = (a+b)(a-b) so, 49x^2 - 81 = (7x)^2 - 9^2 = (7x + 9)(7x - 9)

OpenStudy (raden):

8. the common factor of 6x^3 - 864x is 6x. so, it can be factored. 6x^3 - 864x = 6x (x^2 - 144) = 6x (x^2 - 12^2) = 6x (x+12)(x-12) see one of the factor above is ...

OpenStudy (raden):

9 . the common factor of p^5 - 144p^3 is p^3. it can be written as p^3(p^2 - 144) = p^3 (p^2 - 12^2) = p^3 (p + 12)(p - 12) see, all factore is here

OpenStudy (anonymous):

is it x+12?

OpenStudy (raden):

yes

OpenStudy (anonymous):

okay can you do the ones i didnt answer with work so i can understand them :) please

OpenStudy (raden):

2. 16a^3b^7 + 2a^6b^4 - 22a^4b^5 first, can you determine the common factor of 16, 2, and 22 ? and the common factor of a^3b^7, a^6b^4, a^4b^5 ?

OpenStudy (anonymous):

8 1 11 is it a?